# Text clustering with LLM embeddings

### v3

@@ Abstract
Text clustering is an important approach for organising the growing amount of digital content, helping to structure and find hidden patterns in uncategorised data. However, the effectiveness of text clustering heavily relies on the choice of textual embeddings and clustering algorithms. We argue that recent advances in large language models (LLMs) can potentially improve this task. In this research, we investigated how different textual embeddings—particularly those used in LLMs—and clustering algorithms affect how text datasets are clustered. A series of experiments were conducted to assess how embeddings influence clustering results, the role played by dimensionality reduction through summarisation, and model size adjustment. Findings reveal that LLM embeddings excel at capturing subtleties in structured language, while BERT leads the lightweight options in performance. In addition, we observe that increasing model dimensionality and employing summarization techniques do not consistently lead to improvements in clustering efficiency, suggesting that these strategies require careful analysis to use in real-life models. These results highlight a complex balance between the need for refined text representation and computational feasibility in text clustering applications. This study extends traditional text clustering frameworks by incorporating embeddings from LLMs, providing a path for improved methodologies, while informing new avenues for future research in various types of textual analysis.
@@ 1 Introduction
Text clustering has garnered significant interest in the field of text analysis due to its potential to uncover hidden structures within large sets of unstructured textual data. With the growth of digital text content generated through platforms such as social media, online news outlets, and academic publications, the ability to organise and analyse these data has become increasingly critical. Text clustering can potentially organise large volumes of unstructured data into meaningful categories, enabling efficient information retrieval and insightful thematic analysis across various domains, for example, customer feedback, academic research, and social media content. Text clustering is a preliminary step in other text analysis tasks such as topic modelling, trend analysis, and sentiment analysis. By grouping similar texts, subsequent analyses can proceed with enhanced accuracy and relevance, focusing on more homogeneous data sets with particular characteristics or themes.
@@
As an analytical task, text clustering involves grouping text documents into clusters such that texts within the same cluster are more similar to each other than to those in different clusters. It relies on the principle that text documents can be mathematically represented as vectors in a high-dimensional space, also called embeddings, where dimensions correspond to various features extracted from the documents, such as word frequency or context. Clustering algorithms exploit measures of proximity or resemblance to group documents that exhibit close correspondence in the feature space.
@@
The research presented in this article aims to contribute to the domain of text clustering by testing and identifying optimal combinations of embeddings—including those used in recently released large language models (LLMs)—and clustering algorithms that maximise clustering performance across various datasets. More specifically, the main goal of this paper is to determine if embeddings derived from LLMs outperform traditional embedding techniques, such as Term Frequency-Inverse Document Frequency (TF-IDF). Additionally, we conduct experiments to evaluate the impact of model size and dimensionality reduction through summarisation techniques on clustering performance. Overall, results indicate that LLM embeddings are highly effective at capturing the structured aspects of language, while BERT demonstrates superior performance among lightweight options. Furthermore, increasing model dimensionality and employing summarisation techniques do not consistently enhance clustering efficiency, suggesting that these strategies require careful evaluation for practical application. These findings underscore the need to balance detailed text representation with computational feasibility in text clustering tasks.
@@
This paper is organised as follows. In Section 2, we describe advancements in textual embeddings and briefly mention classical text clustering algorithms used in this domain. In Section 3, we outline the main steps and components of this study, namely dataset selection, data preprocessing, embeddings, and clustering algorithms configuration used to assess clustering quality. In Section 4, we present the results of our study and provide a discussion of these findings. Limitations encountered during the study, along with recommendations for overcoming them, are acknowledged in Section 5. Finally, in Section 6, we synthesise the main conclusions of this research and suggest future developments.
@@ 2.1 Text embeddings
The landscape of text representation in natural language processing (NLP) has undergone an impressive transformation over the past few decades. From simplistic representations to highly sophisticated embeddings, advancements in this domain have played an important role in enabling machines to process and understand human language with increasing accuracy.
One of the earliest methods of text representation that laid the groundwork for subsequent advances was Term Frequency-Inverse Document Frequency (TF-IDF). This method quantifies the importance of a word within a document relative to a corpus, accounting for term frequency and inverse document frequency [1]. While TF-IDF effectively highlighted the relevance of words, it treated each term as independent and failed to capture word context and semantic meaning [2].
Word embeddings, such as those produced by Word2Vec [3] and GloVe [4], marked a significant step forward by generating dense vector representations of words based on their contexts. These models leveraged surrounding words and large corpora to learn word relationships, successfully capturing a range of semantic and syntactic similarities. Despite their effectiveness in capturing semantic regularities, these models still provided a single static vector per word, which posed limitations in capturing polysemous words.
@@
Today, LLMs like OpenAI’s GPT are at the forefront of generating state-of-the-art embeddings [6]. LLMs extend the capabilities of previous models by providing an unprecedented depth and breadth of knowledge encoded in word and sentence-level embeddings. These models are trained on extensive datasets to capture a broad spectrum of human language variations and generate embeddings that reflect a comprehensive understanding of contexts and concepts. The progression from TF-IDF to sophisticated LLM embeddings represents a significant advancement towards more contextually aware text representation in NLP. This evolution continues to propel the field forward, expanding the possibilities for applications like text clustering, sentiment analysis, and beyond.
@@ 2.2 Text clustering algorithms
Text clustering involves grouping a set of texts so that texts in the same group (referred to as a cluster) are more similar to each other than those in different clusters. This section overviews classic clustering algorithms widely used for clustering textual data.
@@
In turn, agglomerative hierarchical clustering (AHC) builds nested clusters by merging them successively. This bottom-up approach starts with each text as a separate cluster and combines clusters based on a linkage criterion, such as the minimum or maximum distance between cluster pairs [8]. AHC is versatile and allows for discovering hierarchies in data, but it can be computationally expensive for large datasets.
@@
Fuzzy $c$-means (FuzzyCM) is a clustering method that allows data points to belong to more than one cluster. This method minimises the objective function with respect to membership and centroid position; it provides soft clustering and can handle overlapping clusters [10]. FuzzyCM is useful when the boundaries between clusters are not clearly defined, although it is computationally more intensive than $k$-means.
In addition to these classic approaches, recent years have seen the rise of alternative methods for text clustering that leverage the strengths of modern embeddings and consider the unique properties of textual data. These include using deep learning models, particularly those based on autoencoders, to learn meaningful low-dimensional representations ideal for clustering [11, 12]. Others explore graph-based clustering methods that construct a graph from text embeddings and partition this graph into clusters [13]. Ensemble clustering approaches, where multiple clustering algorithms are combined to improve the robustness and quality of the results, have also been gaining attention. These methods benefit from the diversity of the individual algorithms and can sometimes overcome the limitations of any single method [14].
The field of text clustering is rapidly expanding, and keeping up with the latest findings is crucial for advancing the state of the art. Recent papers, such as those exploring the use of transformer-based embeddings for clustering [15] or integrating external knowledge bases into the clustering process [16], represent just the tip of the iceberg in this area of research. Additionally, there are some works exploring the potential of using semi-supervised graph clustering for NLP [17]. Moreover, an emerging technique involves using summarisation to refine text clustering, positing that by condensing text into more manageable representations, we can achieve clusters of greater coherence and interpretability—a notion that is particularly advantageous when faced with large and complex datasets [18]. Finally, very recent work by Keraghel and el. [19] offers a preliminary discussion on the potential of using LLM embeddings for text clustering. The authors analysed BLOOMZ, Mistral, Llama-2, and OpenAI embeddings with five clustering algorithms. This paper extends that work by focusing on different datasets, testing additional LLM embeddings, and comparing results using classical TF-IDF as a baseline. Furthermore, this study experiments with summarisation as a dimensionality reduction technique and evaluates the impact of model size on clustering results.
@@ 3 Methods
This research evaluates the effectiveness of various embedding representations in enhancing the performance of text clustering algorithms, aiming to identify the most informative embeddings for clustering tasks. For this purpose, we systematically experimented with multiple datasets, embeddings, and clustering methods and performed an in-depth analysis of clustering results using different evaluation metrics. Overall, the following steps were taken during the course of this study:
@@
    Selection of datasets.
@@ 3.1 Datasets
We selected five datasets to cover a variety of text clustering challenges. Table 1 shows these datasets and their characteristics. The CSTR abstracts dataset [20] is a corpus of 299 scientific abstracts from the Centre for Speech Technology Research. The homogeneous and domain-specific nature of the CSTR dataset allowed us to investigate the effectiveness of clustering techniques in discerning fine-grained topic distinctions in scholarly text related to categories such as artificial intelligence, theory, systems, and robotics. The SyskillWebert dataset [21], which includes user ratings for web pages, enabled exploratory clustering to analyse information used in recommendation systems. The 20Newsgroups dataset [22] is a famous collection of approximately 19,000 news documents partitioned across 20 different classes. With its broad assortment of topics and noisy, unstructured text, this dataset provided a realistic scenario for evaluating the robustness of clustering algorithms under less-than-ideal conditions. To complement these, we added the MN-DS dataset [23], a diverse compilation of multimedia news articles, which provides the opportunity to explore the effectiveness of clustering algorithms in handling multi-categorical data. The dataset is organised hierarchically in two levels. Therefore, experiments were performed independently for each level. Additionally, we included the Reuters dataset [24], a benchmark dataset widely used in text mining and information retrieval research that contains stories from the Reuters news agency, to further enhance the comprehensiveness of our evaluation. We selected documents assigned to single classes to establish a clear ground truth for evaluation, reducing the dataset from 10,369 to 8,654 text articles. This allowed us to assess the clustering algorithms’ ability to accurately group similar documents according to their predefined categories.
@@ 3.2 Preprocessing
For all datasets, a series of preprocessing steps were taken to ensure the quality and uniformity of the input data. The initial step involved removing miscellaneous items such as irrelevant metadata, HTML tags, and any extraneous content that might skew the analysis. Continuing with this methodology, we systematically eliminated invalid non-Latin characters from the dataset. These characters could arise from multilingual data sources or artefacts of data collection, such as encoding errors. Given that the employed workflow is optimised for Latin-based languages, retaining such characters could increase the dimensionality of the feature space, therefore undermining clustering performance, both in terms of computational efficiency and interpretability of the results.
@@ 3.3 Text embeddings
For textual data representation, we compared classical embeddings [5, 26] and state-of-the-art LLM embeddings. Traditional TF-IDF vectors served as a baseline, providing a sparse but interpretable representation based on word importance and analysed dataset. BERT embeddings were extracted from the BERT model, which was trained as a transformer-based bidirectional encoder on BookCorpus [27] and English Wikipedia [28]. These embeddings were utilised to achieve deep contextual understanding, capturing the semantic variations across the corpus. We also employed “text-embedding-ada-002” [29] embeddings since they demonstrated the best results among OpenAI’s embeddings on larger datasets for text search, code search, and sentence similarity tasks. Additionally, other LLM embeddings, computed for Falcon [30] and LLaMA-2-chat [31] models, were included for their respective advancements in performance and efficiency. Falcon embeddings were trained on a hybrid corpus consisting of both text documents and code, while the LLaMA-2-chat embeddings—built on the foundation of LLaMA-2 model using an optimised auto-regressive transformer—underwent targeted fine-tuning for dialogue and question-and-answer tasks. BERT, Falcon, and LLaMA-2-chat embeddings were obtained from Hugging Face’s transformers library, where Hugging Face is a company specialising in providing state-of-the-art models in the form of pipelines and making them accessible through their platform [32]. Table 2 describes the exact embeddings and configurations used.
@@ 3.4 Clustering algorithms
The selected clustering algorithms address the diverse nature of text data, which often contains complex patterns requiring robust methods for effective grouping. Standard $k$-means clustering was used for its simplicity and efficiency in dealing with large datasets. $K$-means++ was chosen as an enhanced variant of $k$-means, with careful initialisation to improve convergence and cluster quality [33]. AHC was utilised for its ability to reveal nested structures within the data. Compared to $k$-means, which assigns each data point to a single cluster, FuzzyCM provides a probabilistic membership approach, accommodating polysemy and subtle semantic differences typical in text data. Finally, Spectral clustering was selected for its effectiveness in identifying clusters based on their data-induced graph structure and also since it is particularly adept at discovering clusters with non-convex shapes. We used Euclidean distance as the document similarity function and employed the nearest centroid approach to map clusters to original labels for all clustering algorithms.
The selected algorithms and their respective parameters are listed in Table 3. The scikit-learn library [34] provided the implementations for all algorithms except for FuzzyCM, which was used from the scikit-fuzzy package [35].

### v4

@@ Abstract
Text clustering is an important method for organising the increasing volume of digital content, aiding in the structuring and discovery of hidden patterns in uncategorised data. The effectiveness of text clustering largely depends on the selection of textual embeddings and clustering algorithms. This study argues that recent advancements in large language models (LLMs) have the potential to enhance this task. The research investigates how different textual embeddings, particularly those utilised in LLMs, and various clustering algorithms influence the clustering of text datasets. A series of experiments were conducted to evaluate the impact of embeddings on clustering results, the role of dimensionality reduction through summarisation, and the adjustment of model size. The findings indicate that LLM embeddings are superior at capturing subtleties in structured language. OpenAI’s GPT-3.5 Turbo model yields better results in three out of five clustering metrics across most tested datasets. Most LLM embeddings show improvements in cluster purity and provide a more informative silhouette score, reflecting a refined structural understanding of text data compared to traditional methods. Among the more lightweight models, BERT demonstrates leading performance. Additionally, it was observed that increasing model dimensionality and employing summarisation techniques do not consistently enhance clustering efficiency, suggesting that these strategies require careful consideration for practical application. These results highlight a complex balance between the need for refined text representation and computational feasibility in text clustering applications. This study extends traditional text clustering frameworks by integrating embeddings from LLMs, offering improved methodologies and suggesting new avenues for future research in various types of textual analysis.
@@ 1 Introduction
Text clustering has garnered significant interest in the field of text analysis due to its potential to uncover hidden structures within large sets of unstructured textual data. With the exponential growth of digital text content generated through platforms such as social media, online news outlets, and academic publications, the ability to organise and analyse these data has become increasingly critical. Text clustering can organise large volumes of unstructured data into meaningful categories, enabling efficient information retrieval and insightful thematic analysis across various domains, such as customer feedback, academic research, and social media content.
Text clustering serves as a preliminary step in various text analysis tasks, including topic modelling, trend analysis, and sentiment analysis. By grouping similar texts, subsequent analyses can proceed with enhanced accuracy and relevance, focusing on more homogeneous data sets with specific characteristics or themes. This process improves the precision of the analyses and ensures that the insights derived are more targeted and applicable to the context in which the data were generated.
@@
As an analytical task, text clustering involves grouping text documents into clusters such that texts within the same cluster are more similar to each other than to those in different clusters. This process relies on the principle that text documents can be mathematically represented as vectors in a high-dimensional space, known as embeddings, where dimensions correspond to various features extracted from the documents, such as word frequency or context. Clustering algorithms use measures of proximity or resemblance to group documents that exhibit close correspondence in the feature space. This approach enables the identification of natural groupings within the data, facilitating more effective organisation and analysis of large text corpora.
@@
The research presented in this article aims to contribute to the domain of text clustering by testing and identifying optimal combinations of embeddings—including those used in recently released large language models (LLMs)—and clustering algorithms that maximise clustering performance across various datasets. The primary objective of this paper is to determine if embeddings derived from LLMs outperform traditional embedding techniques, such as Term Frequency-Inverse Document Frequency (TF-IDF). Additionally, experiments are conducted to evaluate the impact of model size and dimensionality reduction through summarisation techniques on clustering performance.
Results indicate that LLM embeddings are highly effective at capturing the structured aspects of language, with BERT demonstrating superior performance among lightweight models. Moreover, increasing model dimensionality and employing summarisation techniques do not consistently enhance clustering efficiency, suggesting that these strategies require careful evaluation for practical application. These findings underscore the need to balance detailed text representation with computational feasibility in text clustering tasks.
@@
This paper is organised as follows. In Section 2, advancements in textual embeddings are described, and classical text clustering algorithms used in this domain are briefly mentioned. Section 3 outlines the main steps and components of this study, including dataset selection, data preprocessing, embeddings, and clustering algorithm configurations used to assess clustering quality. Section 4 presents the results of our study and provides a discussion of these findings. Limitations encountered during the study, along with recommendations for overcoming them, are acknowledged in Section 5. Finally, Section 6 synthesises the main conclusions of this research and suggests future developments.
@@ 2.1 Text Embeddings
The landscape of text representation in natural language processing (NLP) has undergone an impressive transformation over the past few decades. From simplistic representations to highly sophisticated embeddings, advancements in this domain have significantly enhanced the ability of machines to process and understand human language with increasing accuracy.
One of the earliest methods of text representation that laid the groundwork for subsequent advances was Term Frequency-Inverse Document Frequency (TF-IDF). This method quantifies the importance of a word within a document relative to a corpus by accounting for term frequency and inverse document frequency [1]. While TF-IDF effectively highlights the relevance of words, it treats each term as independent and fails to capture word context and semantic meaning [2].
Word embeddings, such as those produced by Word2Vec [3] and GloVe [4], marked a significant advancement by generating dense vector representations of words based on their contexts. These models leveraged surrounding words and large corpora to learn word relationships, successfully capturing a range of semantic and syntactic similarities. Despite their effectiveness in capturing semantic regularities, these models still provided a single static vector per word, which posed limitations in handling polysemous words—words with multiple meanings.
@@
Today, LLMs like OpenAI’s GPT are at the forefront of generating state-of-the-art embeddings [6]. LLMs extend the capabilities of previous models by providing an unprecedented depth and breadth of knowledge encoded in word and sentence-level embeddings. These models are trained on extensive datasets to capture a broad spectrum of human language variations and generate embeddings that reflect a comprehensive understanding of contexts and concepts.
The progression from TF-IDF to sophisticated LLM embeddings represents a significant advancement towards more contextually aware text representation in NLP. This evolution continues to propel the field forward, expanding the possibilities for applications such as text clustering, sentiment analysis, and beyond. In the context of text clustering methodologies, there exists a considerable research gap that underscores the need for a comprehensive evaluation of LLM embeddings against traditional techniques such as TF-IDF.
@@ 2.2 Text Clustering Algorithms
Text clustering involves grouping a set of texts such that texts in the same group (referred to as a cluster) are more similar to each other than to those in different clusters. This section provides an overview of classic clustering algorithms widely used for clustering textual data.
@@
Agglomerative hierarchical clustering (AHC) builds nested clusters by merging them successively. This bottom-up approach starts with each text as a separate cluster and combines clusters based on a linkage criterion, such as the minimum or maximum distance between cluster pairs [8]. AHC is versatile and allows for discovering hierarchies in data, but it can be computationally expensive for large datasets.
@@
Fuzzy $c$-means (FuzzyCM) is a clustering method that allows data points to belong to more than one cluster. This method minimises the objective function with respect to membership and centroid position, providing soft clustering and handling overlapping clusters [10]. FuzzyCM is useful when the boundaries between clusters are not clearly defined, although it is computationally more intensive than $k$-means.
In addition to these classic approaches, recent years have seen the rise of alternative methods for text clustering that leverage the strengths of modern embeddings and consider the unique properties of textual data. These include using deep learning models, particularly those based on autoencoders, to learn meaningful low-dimensional representations ideal for clustering [11, 12]. Ensemble clustering approaches, where multiple clustering algorithms are combined to improve the robustness and quality of the results, have also been gaining attention. These methods benefit from the diversity of the individual algorithms and can sometimes overcome the limitations of any single method [13].
The field of text clustering is rapidly expanding, and keeping up with the latest findings is crucial for advancing the state of the art. Recent papers, such as those exploring the use of transformer-based embeddings for clustering [14] and integrating external knowledge bases into the clustering process [15], represent just the tip of the iceberg in this area of research. Additionally, very recent work by Keraghel et al. [16] offers a preliminary discussion on the potential of using LLM embeddings for text clustering. The authors analysed embeddings from BLOOMZ, Mistral, Llama-2, and OpenAI using five clustering algorithms. This paper extends that work by focusing on different datasets, testing additional LLM embeddings, and comparing results using classical TF-IDF as a baseline. Furthermore, this study experiments with summarisation as a dimensionality reduction technique and evaluates the impact of model size on clustering results.
@@ 3 Methods
This research evaluates the effectiveness of various embedding representations in enhancing the performance of text clustering algorithms, aiming to identify the most informative embeddings for clustering tasks. To achieve this, we systematically experimented with multiple datasets, embeddings, and clustering methods, and performed an in-depth analysis of clustering results using different evaluation metrics. The following steps were undertaken during the course of this study:
@@
    Selection of datasets, ensuring the robustness of our findings across different types of textual data.
@@ 3.1 Datasets
We selected five datasets to cover a variety of text clustering challenges. Table 1 shows these datasets and their characteristics. The CSTR abstracts dataset [17] is a corpus of 299 scientific abstracts from the Centre for Speech Technology Research. The homogeneous and domain-specific nature of the CSTR dataset allowed us to investigate the effectiveness of clustering techniques in discerning fine-grained topic distinctions in scholarly text related to categories such as artificial intelligence, theory, systems, and robotics.
The SyskillWebert dataset [18], which includes user ratings for web pages, enabled exploratory clustering to analyse information used in recommendation systems. The 20Newsgroups dataset [19] is a famous collection of approximately 19,000 news documents partitioned across 20 different classes. With its broad assortment of topics and noisy, unstructured text, this dataset provided a realistic scenario for evaluating the robustness of clustering algorithms under less-than-ideal conditions.
To complement these, we added the MN-DS dataset [20], a diverse compilation of multimedia news articles, which provides the opportunity to explore the effectiveness of clustering algorithms in handling multi-categorical data. The dataset is organised hierarchically in two levels; therefore, experiments were performed independently for each level. Additionally, we included the Reuters dataset [21], a benchmark dataset widely used in text mining and information retrieval research, which contains stories from the Reuters news agency. To enhance the comprehensiveness of our evaluation, we selected documents assigned to single classes, reducing the dataset from 10,369 to 8,654 text articles. This allowed us to assess the clustering algorithms’ ability to accurately group similar documents according to their predefined categories.
@@ 3.2 Preprocessing
For all datasets, a series of preprocessing steps were taken to ensure the quality and uniformity of the input data. The initial step involved removing miscellaneous items such as irrelevant metadata, HTML tags, and any extraneous content that might skew the analysis. Continuing with this methodology, we systematically eliminated invalid non-Latin characters from the dataset. These characters could arise from multilingual data sources or artefacts of data collection, such as encoding errors. Given that the employed workflow is optimised for Latin-based languages, retaining such characters could increase the dimensionality of the feature space, thereby undermining clustering performance, both in terms of computational efficiency and interpretability of the results.
@@ 3.3 Text Embeddings
For textual data representation, we compared classical embeddings [5, 23] and state-of-the-art LLM embeddings. Traditional TF-IDF vectors served as a baseline, providing a sparse but interpretable representation based on word importance within the analysed dataset. BERT embeddings were extracted from the BERT model, which was trained as a transformer-based bidirectional encoder on BookCorpus [24] and English Wikipedia [25]. These embeddings were utilised to achieve deep contextual understanding, capturing the semantic variations across the corpus.
We also employed “text-embedding-ada-002” [26] embeddings, as they demonstrated the best results among OpenAI’s embeddings on larger datasets for tasks such as text search, code search, and sentence similarity. Additionally, other LLM embeddings were computed for Falcon [27] and LLaMA-2-chat [28] models, included for their respective advancements in performance and efficiency. Falcon embeddings were trained on a hybrid corpus consisting of both text documents and code, while the LLaMA-2-chat embeddings—built on the foundation of the LLaMA-2 model using an optimised auto-regressive transformer—underwent targeted fine-tuning for dialogue and question-and-answer tasks.
BERT, Falcon, and LLaMA-2-chat embeddings were obtained from Hugging Face’s transformers library, a platform providing state-of-the-art models in accessible pipelines [29]. Table 2 describes the exact embeddings and configurations used.
@@ 3.4 Clustering Algorithms
The selected clustering algorithms address the diverse nature of text data, which often contains complex patterns requiring robust methods for effective grouping. Standard $k$-means clustering was used for its simplicity and efficiency in dealing with large datasets. $K$-means++ was chosen as an enhanced variant of $k$-means, with careful initialisation to improve convergence and cluster quality [30]. AHC was utilised for its ability to reveal nested structures within the data. Compared to $k$-means, which assigns each data point to a single cluster, FuzzyCM provides a probabilistic membership approach, accommodating polysemy and subtle semantic differences typical in text data. Finally, Spectral clustering was selected for its effectiveness in identifying clusters based on their data-induced graph structure, and it is particularly adept at discovering clusters with non-convex shapes.
To map the newly formed clusters to the original labels of the dataset, we calculated the Euclidean distance between the centroid of each derived cluster and the centroid of each ground truth cluster, assigning the closest ones. The Euclidean distance is a preferable metric for this process because it accurately reflects distances in multidimensional space, ensuring precise mapping. This method allows the use of external evaluation metrics such as the $F1$-score. In our work, we did not apply Euclidean distance at the word level, which can be problematic. Instead, the process computes an aggregation for each identified cluster, enabling a more reliable association with the existing labels.
The selected algorithms and their respective parameters are listed in Table 3. The scikit-learn library [31] provided the implementations for all algorithms except for FuzzyCM, which was utilised from the scikit-fuzzy package [32].